OpenStudy (anonymous):

How can x2 = 2x + 4 be set up as a system of equations? y = x2 - 2x + 4 y = x2 + 2x + 4 y = x2 y = x2 + 2x + 4 y = x2 + 2x y = x2 + 4 y = x2 y = 2x + 4 @ganeshie8 PLeaaase help me out?!

Directrix (directrix):

x^2 = 2x + 4 looks like a step in solving a system of equations by substitution. There are no y variables here so the substitution must or might have been for y. y = x^2 and y = 2x + 4 could be the system of equations. See what you think. @iamabarbiegirl

Directrix (directrix):

There may be more than one correct answer option.

OpenStudy (anonymous):

okay so u think that its D? @Directrix

Directrix (directrix):

Option D is correct but also consider the others.

OpenStudy (anonymous):

oh okay! i get it thanks!! :D do you mind keeping an eye for more questions from me if i get stuck on some?

Directrix (directrix):

Looking at option A y = x2 - 2x + 4, y = x2 + 2x + 4 Substituting for y, x^2 - 2x + 4 = x^2 + 2x + 4 The x^2 terms subtract off so option A is not correct.

Directrix (directrix):

That holds true for all the other options so I would go with option D. ------- Yes, I'll keep an eye out for your problems. You can tag me if you like and I'll see if I know how to help. @iamabarbiegirl

OpenStudy (anonymous):

alright thank you very much! :D I appreciate your help and time

Directrix (directrix):

You are welcome.

OpenStudy (anonymous):

please check this for me @Directrix Brad sold candy bars and cookies for a fundraiser at school. Candy bars sold for $3, and cookies sold for $5. He sold a total of 20 items and made $76. How many candy bars did Brad sell? 12<----- MY ANSWER. 14 16 18

Directrix (directrix):

I had to crank out the equations and solve. I got 8 cookies and 12 bars.

OpenStudy (anonymous):

yayyy!! I got it right :D thanks @Directrix

Directrix (directrix):

Good to know.

OpenStudy (anonymous):

How can x + 4 = 3x + 12 be set up as a system of equations? y = x + 1 y = -3x + 12 y = x + 2 y = -3x + 12 y = x + 4 <------- MY ANSWER y = 3x + 12 y = 4 y = -3x + 12 @Directrix

OpenStudy (anonymous):

stuck on this one @Directrix Which equation does the graph of the systems of equations solve? quadratic graph opening down and quadratic graph opening up. They intersect at 0, 3 and 3, 0. x2 - 2x + 3 = 2x2 - 4x - 3 x2 - 2x + 3 = 2x2 - 4x + 3 -x2 + 2x + 3 = x2 - 4x + 3 -x2 - 2x + 3 = x2 - 4x + 3

OpenStudy (anonymous):

@ganeshie8 @iambatman please help me out guys.

Directrix (directrix):

On this problem, I agree with your answer. @iamabarbiegirl How can x + 4 = 3x + 12 be set up as a system of equations?